A rectangle has sides of 3x – 4 and 7x + 10. Find the expression that represents its perimeter.
Answer by venugopalramana(3286)

L=3X-4
B=7X+10
PERIMETER=P=2*(L+B)=2(3X-4+7X+10)=2(10X+6)=20X+12
P=20X+12